# Low-Complexity Monitoring and Compensation of Transceiver IQ Imbalance by Multi-dimensional Architecture for Dual-Polarization 16 Quadrature Amplitude Modulation

### v1

@@
††journal: opticajournal††articletype: Research ArticleIn this paper, a low-complexity IQ imbalance compensation architecture is proposed, which reduces the effects of in-phase (I) and quadrature (Q) imbalance. The architecture consists of transceiver IQ skew estimation methods and a low-complexity MIMO equalizer structure. Before the IQ skew estimation, the chromatic dispersion(CD) is pre-compensated in the transmitter(TX) by chirp filtering. The receiver(RX) IQ skew is estimated by Gardner’s phase detector, and the TX skew is estimated by finding the value that yields the lowest equalizer error. The low-complexity MIMO equalizer consists of a complex-valued MIMO(CV-MIMO) and a real-valued DD-LMS MIMO(RV-MIMO), which employ a butterfly and a non-butterfly structure, respectively. The CV-MIMO is used to perform polarization demultiplexing. The RV-MIMO equalizes each of the two polarisations and simultaneously compensates for the TX IQ imbalance. The architecture first compensates for the IQ skew at low-complexity, and the other imperfections are compensated by the low-complexity MIMO equalizer. Therefore, this architecture can equalize signals impaired by the transceiver IQ imbalance with low complexity. A 100 km transmission simulation and experiment with 36 Gbaud dual-polarization quadrature amplitude modulation(DP-QAM) signals and offline DSP showed that, with the RX IQ skew estimation, the number of real multiplications is reduced by more than 70% compared with conventional cases. With the low-complexity MIMO equalizer, the number of real multiplications is reduced by 51% compared with 4×4 MIMO.
@@ 1 Introduction
To accommodate the increasing demand for data rate, passive optical network(PON) are transitioning to high symbol rates and high spectral efficiency modulation formats. High-order quadrature amplitude modulation (QAM) with a high baud rate has the potential to increase the data rate and efficiency. However, the QAM signal require more complex transceiver to send and receive it. For example, the dual-polarisation optical QAM signal must pass through a dual-polarization optical hybrid, balanced photodetector(BPD) and transimpedance amplifier(TIA) to be demodulated. These components may impair the signal under unideal conditions. In an unideal condition, it is common to have a timing misalignment (IQ skew) of several picoseconds between the in-phase(I) and quadrature (Q)paths, or amplitude/phase imbalance between I and Q. Such as waveguide length mismatches between I/Q arms, TIA delay and gain imbalance, or any imperfection in the optical front-end. These imbalances are commonly calibrated during the manufacturing process. However, some residual imbalances can remain. In addition, environmental and mechanical factors, including temperature fluctuations and device aging, cause dynamic changes in IQ imperfection. At higher symbol rates, this post-calibration residual imbalance has a more severe impact than on lower symbol-rate transmissions. The technical document for the 400ZR coherent optical interface[30] stipulates that IQ skew should be less than 0.75 ps while IQ amplitude and phase imbalances must not exceed 1 dB and ±5° respectively. Thus, for high symbol-rate transmissions, an algorithm robust to IQ imbalance or a method to estimate the IQ imbalance is required.
@@
Recently, some IQ imbalance-tolerant MIMO adaptive equalizers were proposed. References [1] propose equalizing to each in-phase and quadrature in each polarization, having four real-valued inputs and four real-valued outputs. This is a well-known real-valued 44 MIMO structure. References [2] propose independent chromatic dispersion (CD) compensation for each I and Q. After that, References [2] perform a 42 MIMO adaptive equalizer to compensate for IQ imbalance. References [3] propose an 82 MIMO adaptive equalizer structure. The new structure not only compensates for the IQ imbalance in single polarization but also for the presence of arbitrary crosstalk across four IQ lanes of polarization-multiplexed transceivers. References [4, 5] propose a complex-valued adaptive equalizer structure that can compensate for linear IQ-mixing effects at the receiver(RX) front end. References [6] proposed using the strictly linear and widely linear filters to compensate for the IQ imbalance over 10000 km. However, the conventional receiver IQ imbalance compensation methods based on the AEQ suffer from IQ skew-enhanced timing jitter incurred by the clock recovery algorithm[9]. In addition, imbalance-tolerant MIMO equalizers require more taps and multiple iterations, which introduces an increase in complexity. Therefore, skew-tolerant MIMO demands more area in practical ASIC implementations in comparison to common complex-valued 2×2 MIMO equalizers.
@@
Therefore, some IQ skew monitor algorithms based on a clock recovery algorithm were proposed. References [7, 8, 9] use a digital Godard’s phase detector (Godard’s PD) to monitor the transceiver’s IQ skew. References [10, 11] propose a robust receiver IQ skew monitoring scheme using a modified Godard’s PD algorithm tailored for non-integer oversampling receivers. References [12, 13, 14] propose a receiver IQ skew monitoring scheme using a modified Gardner’s phase detector (Gardner’s PD) algorithm or an absolute phase detector algorithm. In general, Gardner’s PD is better than Godard’s PD because Gardner’s PD doesn’t require the extra FFT. However, the IQ skew monitor based on Gardner’s PD algorithm is suitable only for short fiber transmission length due to fiber polarization mode dispersion (PMD) and chromatic dispersion (CD). When the signal polarization rotation (SOP) angle crosses 42.5∘ and the differential group delay (DGD) is 1/2 symbol period [7, 8], the algorithms will fail in Gardner’s PD algorithms [15]. The skew monitor using Gardner’s PD is influenced by CD compensation, which will be studied in subsequent content. Some methods estimate the IQ imbalance based on data-aid. Reference [16] uses the concatenation of the pure real and imaginary sequences in the time domain to estimate the transceiver IQ skew. Reference[17] proposed a scheme based on cross-power-spectrum analysis for IQ imperfections. The proposed cross-power-spectrum scheme exhibits up to ±200 ps for the skew of the coherent optical transceivers. To achieve wide-range IQ skew calibration simultaneously, Reference[18] studies propose utilizing the cross-correlation of time domain signals to extract delay information.
@@
Because the IQ imbalance varies slowly, it is unnecessary to monitor it frequently. The current methods that compensate for the IQ imbalance are inefficient. For example, equalizers add more taps to compensate for receiver IQ imbalance. The methods that estimate the IQ imbalance based on data-aid need extra area in the circuit and increase the complexity. In addition, the IQ skew is the primary factor influencing the bit error rate(BER). Reference[19] clearly shows that the effective SNR reduction of the three factors is skew>amp. imbalance>pha. imbalance. Reference[20] shows that for a 100 GBaud 16QAM signal, the IQ skew should be less than 1.1 ps when the signal-to-noise ratio(SNR) penalty is less than 1 dB. Therefore, the IQ skew should first be extracted using low-hardware-resource methods, and then MIMO can reduce the number of taps to euqalize signal. Skew monitoring based on Godard’s PD is a low-hardware-resource method, but it needs FFTs.
@@
In this paper, we introduce a low-complexity IQ imbalance compensation architecture for DP-16QAM in 100 km single-mode fiber(SSMF) transmission. First, the transceiver IQ skew is estimated and compensated. We use Gardner’s PD to estimate the receiver IQ skew. Due to the low PMD coefficient, the DGD over 100 km of SSMF does not approach half the symbol period. Therefore, Gardner’s PD remains robust even when the SOP angle crosses 42.5∘, since the required concurrent condition of DGD = T/2 is not satisfied. Rx IQ skew is added to a signal after accumulation of CD, but when CD compensation is performed first, the IQ impairments become mixed in a complicated way. Thus, we set the CD compensation in the transmitter(TX) DSP to avoid this problem. Moreover, for low complexity, we select the chirp filtering technique to compensate CD. Then, we configure a scanning method to estimate the TX IQ skew. The scanning method estimates and compensates the TX IQ skew by finding the value that yields the lowest equalizer error. Second, we designed a low-complexity equalizer to compensate for the TX amplitude and phase imbalance. We combine the complex-valued MIMO(CV-MIMO) and the real-valued MIMO(RV-MIMO) for signal equalization, which employ a butterfly and a non-butterfly structure, respectively. The CV-MIMO is used to perform polarization demultiplexing. The RV-MIMO equalizes each of the two polarisations and simultaneously compensates for the TX IQ imbalance. Since the TX and RX skew are being compensated, the new equalizer can equalize the signal with fewer taps. For convenience, the new equalizer is called C-R AEQ. The simulation results demonstrate that the RX skew estimation error is less than 0.2 ps, and the TX IQ skew estimation error is less than one-sixteenth of a symbol period. The residue TX IQ skew can be compensated by C-R AEQ. The C-R AEQ tolerates phase imbalance up to ±10 degrees and amplitude imbalance up to ±0.15. Compared to the 4×4 MIMO equalizer, the C-R AEQ reduces the complexity by approximately half. In the experiment, we conducted 100 km SSMF transmissions following in-field calibration using the proposed method. The study utilized 36 GBaud 16 QAM signals to investigate the enhancement of Q-factor under different IQ balance conditions. The experimental results corroborate the simulation conclusion as well.
@@ 2.1 Working principle
Where the , and . According to Eq. 2, the delay introduces respective linear phases shift to and in frequency domain. Therefore, the transmission matrix of delay in the frequency domain can be calculated by
@@
However, if there are IQ skews in X polarization or Y polarization, the numberically controlled oscillator(NCO) outputs leading to the error sampling instant, as shown in Fig. 1(b). So, if we calculate the I and Q separately, as shown in Fig.1(c), then we can calculate the value of IQ skew by
Where are the receiver IQ skew of the X polarization and Y polarization(Pol), Baud is the symbol rate. 2Baud is the sample rate. The notation mean() represents the average calculation. are fractional intervals of I and Q in each Pol.
However, the CD compensation will affect the algorithm. If the fiber has a CD and the CD has been compensated in the receiver, the skew monitoring cannot calculate the receiver IQ skew. The reason is explained using the WL mode [4]. Let ,we can write . Thus, the , which suffers from CD and receiver IQ skew, can be represented by:
@@
Where is the signal after CD compensation in the frequency domain. Skew monitoring requires extracting the phase shift from . The frequency domain of correspond to . However, from Eq.11, Eq.2, and Eq.3, it can be found that the can not be derived from . Because contains a component which consists of the complex conjugate of the . In other words, if , according to Eq.11, the . Therefore, the receiver IQ skew cannot be detected when the CD is compensated in the receiver. Fig.2 shows this difference. Fig.2(a) shows the phase shift between the delayed signal and the original signal in the frequency domain. It is observed in Fig.2(a) that the delayed signal always keeps a fixed phase difference between the delayed signal and the original signal. This means that the delay can be detected by the skew monitoring. Fig.2(b) shows the phase shift between the and the original signal in the frequency domain. It is observed in Fig.2(b) that the phase difference changes when the changes. According to Eq.3, the delay cannot be detected by the skew monitoring.
To further illustrate the CD compensation and IQ skew interaction at the receiver. Fig.3(a) is the skew monitoring, which is set at the receiver. The TED is Gardner’s PD. Xi, Xq, Yi, and Yq are the I and Q of X polarization and Y polarization, respectively. The skew calculation module is composed of four CRA. Xi, Xq, Yi, and Yq input the four CRA, respectively, and each CRA outputs a fractional interval to obtain the , then, the IQ skew for the corresponding polarization is obtained using Eq.4 and Eq.5. The CD compensation is set before skew monitoring. In Fig. 3(b), the estimation results are plotted as a function of fiber length, for a CD coefficient= in fiber. It should be noted that the skew monitoring runs after CD compensation is completed when the transmission length is greater than 10 km. And the transmission length is less than 10 km, the CD compensation is not performed. The RX IQ skew is set to 3.9 ps in the receiver. After the CD compensation is complete, the skew calculation module computes the skews. According to Fig. 3(b), the skew estimation is close to 3.9 ps when the transmission length is less than 10 km, and the skew estimate algorithm will fail when the transmission length is greater than 10 km. In general, receiver IQ skew is uncorrelated with the transmission length. However, from Fig. 3(b), if the optical signal has a larger CD and the CD is compensated in the receiver, the receiver IQ skew cannot be detected via our proposed skew monitoring.
In fact, the RX IQ implitude and phase imbalance are similar to the RX skew. The RX amplitude and phase imbalance can be written as
Where g is the amplitude imbalance, and is the phase imbalance.If the signal suffers from CD, which can be written as:
Similar to RX skew, the is used to compensate, and the result can be written as
Similar to EQ.11, contains a component which consists of the complex conjugate of the signal. Therefore, IQ amp./phase compensation should be set before CD compensation when CD compensation is in the receiver.
Based on the analysis above, we design a DSP flow as shown in Fig. 4. In TX DSP, to solve the problem that the receiver’s IQ skew cannot be detected with a large CD. We set the CD compensation algorithm at the TX DSP. Furthermore, to reduce the complexity of CD compensation, we select the chirp filter technique. The pre-chirp filter can be written as[22]
@@
The scanning module is used to roughly compensate for the TX IQ skew. The Fig. 5 is the diagram of the scanning module. The scanning module includes a scanning delay, a scan controller, and a rough TX IQ skew compensation. In scanning delay, the X pol, Y pol introduce a gradually increasing delay from -3T/16 to 3T/16, which can be expressed as
Where the and are the delays of X pol and Y pol. The and are a gradually increasing delay from -3T/16 to 3T/16. The RX DSP CV-MIMO outputs the X pol error and Y pol equalize error after each performs Eq.17 and Eq. 18, which can be expressed as
Where N is the number of symbols. Then the X pol and Y pol are sent to the scan controller. It should be noted that the and are respective scanning from -3/16T to 3/16T, the scan controller stores a number of ) (=(-3T/16, -2T/16…2T/16,3T/16)) and ) (=(-3T/16, -2T/16…2T/16,3T/16)). The scan controller selects the minimum M(e(n)x) and M(e(n)y). Then, the and are the rough skew estimation, and the rough TX IQ skew compensation module uses and to compensate for the TX IQ skew. The RX skew monitor is shown in Fig. 3(a).
To fine-compensate for the TX IQ skew, inspired by KM AEQ[23], we design a C-R AEQ, which is composed of complex-valued MIMO(CV-MIMO) and real-valued decision-directed least-mean-square (RV-DD-LMS). The CV-MIMO is composed of a butterfly structure with 6 taps per filter. The RV-DD-LMS is a non-butterfly filter for each of two polarizations after the FOE/CPR. Because we found that the FOE/CPE can be correctly performed after the signal is equalised by only a 3-tap CV-MIMO. The CV-MIMO is used for polarisation demultiplexing and to ensure the FOE/CPE can be correctly performed, and the RV-DD-LMS compensates the TX IQ imbalance and equalises each of the two polarisations. The tap coefficient vector of the RV-DD-LMS is updated according to reference [1]. C-R AEQ is similar to reference[24]. The difference is that the reference[24] does not consider CD and frequency offset. This means reference[24] is only suitable for self-homodyne coherent optical system in short transmission length.

### v2

@@
††journal: opticajournal††articletype: Research ArticleIn this paper, a low-complexity multi-dimensional architecture for IQ imbalance compensation is proposed, which reduces the effects of in-phase (I) and quadrature (Q) imbalance. The architecture use a transceiver IQ skew estimation structure to compensate for IQ skew, and then use a low-complexity MIMO equalizer to compensate for IQ amplitude/phase imbalance. In the transceiver IQ skew estimation structure, the receiver(RX) IQ skew is estimated by Gardner’s phase detector, and the transmitter TX skew is estimated by finding the value that yields the lowest equalizer error. The low-complexity MIMO equalizer consists of a complex-valued MIMO (CV-MIMO) and a two-layer multimodulus algorithm real-valued MIMO (TMMA-RV-MIMO), which employ a butterfly and a non-butterfly structure, respectively. The CV-MIMO is used to perform polarization demultiplexing and the TMMA-RV-MIMO equalizes each of the two polarizations. In addition, the TMMA-RV-MIMO can recovery the carrier phase. A 100 km transmission simulation and experiment with 36 Gbaud dual-polarization 16 quadrature amplitude modulation (DP-16QAM) signals showed that, with the TX/RX IQ skew estimation, the estimation error is less than 0.9/0.25 ps. The low-complexity MIMO equalizer can tolerate 0.1 TX IQ amplitude imbalance and 5 degrees at a 0.3 dB Q-factor penalty. The number of real multiplications is reduced by 55% compared with conventional cases in total.
@@ 1 Introduction
To accommodate the increasing demand for data rates, optical communication systems are transitioning to higher symbol rates and high-spectral-efficiency modulation formats. High-order quadrature amplitude modulation (QAM) with a high baud rate has the potential to increase the data rate and spectral efficiency. However, the QAM signal requires a complex transceiver to send and receive. For example, the dual-polarization optical QAM signal passes through a dual-polarization optical hybrid, a balanced photodetector(BPD) and a transimpedance amplifier(TIA) to be demodulated. These components may impair the signal under non-ideal conditions. In a non-ideal condition, it is common to have a timing misalignment (IQ skew) of several picoseconds between the in-phase (I) and quadrature (Q) paths, or amplitude/phase imbalance between I and Q. These impairments are caused by waveguide length mismatches between I/Q arms, TIA delay and gain imbalance, or other imperfections in the optical front-end. These imbalances are commonly calibrated during the manufacturing process. However, some residual imbalances can remain. In addition, environmental and mechanical factors, including temperature fluctuations and device aging, cause dynamic changes in IQ imperfection. At higher symbol rates, this post-calibration residual imbalance has a more severe impact than on lower symbol-rate transmissions. The technical document for the 400ZR coherent optical interface stipulates that IQ skew should be less than 0.75 ps while IQ amplitude and phase imbalances must not exceed 1 dB and ±5° respectively. Thus, for high symbol-rate transmissions, an algorithm robust to IQ imbalance or a method to estimate the IQ imbalance is required.
@@
Recently, several IQ imbalance-tolerant MIMO adaptive equalizers have been proposed. A well-known real-valued 44 MIMO structure[1] was proposed to compensate for IQ imbalance. However, the structure will fail when the CD is compensated for using a static equalizer[2]. To overcome this problem, some new MIMO structures [2, 3, 4, 5, 6] were proposed to compensate for IQ imbalance. However, the conventional receiver IQ imbalance compensation methods based on the AEQ suffer from IQ skew-enhanced timing jitter incurred by the clock recovery algorithm[9]. In addition, imbalance-tolerant MIMO equalizers require more taps and multiple iterations, which introduces an increase in complexity. Skew-tolerant MIMO demands more area in practical ASIC implementations in comparison to common complex-valued 2×2 MIMO equalizers.
@@
Therefore, some IQ skew monitor algorithms based on a clock recovery algorithm were proposed. Reference [7, 8, 9, 10, 11] use a digital Godard’s phase detector (Godard’s PD) to monitor the transceiver’s IQ skew. Reference [12, 13, 14] propose receiver IQ skew tolarent scheme using a modified Gardner’s phase detector (Gardner’s PD) algorithm. In general, Gardner’s PD is better than Godard’s PD because Gardner’s PD doesn’t require the extra FFT. However, the IQ skew monitor based on Gardner’s PD algorithm is suitable only for short fiber transmission length due to fiber polarization mode dispersion (PMD) and chromatic dispersion (CD) compensation algorithm. When the signal polarization rotation (SOP) angle crosses 42.5∘ and the differential group delay (DGD) is 1/2 symbol period, the skew estimation will fail in Gardner’s PD algorithms [7, 10, 15]. Furthermore, the skew monitor using Gardner’s PD is influenced by CD compensation. The reason is thar Rx IQ skew are added to a signal after accumulation of CD, but when CD compensation is performed first, the IQ skew become mixed in a complicated way. Some methods estimate the IQ imbalance based on data-aid. Reference [16, 17, 18] are able to achieve wide-range IQ skew calibration simultaneously based training symbols.
@@
Because the IQ imbalance varies slowly, it is unnecessary to monitor it frequently. The current methods that compensate for the IQ imbalance are inefficient. For example, equalizers add more taps to compensate for receiver IQ imbalance. The methods that estimate the IQ imbalance based on data-aid need extra area in the circuit and increase the complexity. In addition, the IQ skew is the primary factor influencing the Q-factor in IQ imbalance. Reference.[19] shows that the effective SNR reduction factor is IQ skew. Reference[20] shows that for a 100 GBaud 16QAM signal, the IQ skew should be less than 1.1 ps when the signal-to-noise ratio(SNR) penalty is less than 1 dB. Therefore, the IQ skew should first be extracted using low-hardware-resource methods, and then equalizer can reduce the number of taps to equalize the signal. Skew monitoring based on Godard’s PD is a low-hardware-resource method, but it needs FFTs.
@@
In this paper, we introduce a low-complexity IQ imbalance compensation architecture for DP-16QAM in 100 km single-mode fiber(SSMF) transmission. First, the transceiver IQ skew is estimated and compensated. We use Gardner’s PD to estimate the receiver IQ skew. Due to the low PMD coefficient, Gardner’s PD remains robust in 100 km transmission length. To avoid the IQ mixed, we set the CD compensation in the transmitter(TX) DSP. Moreover, for low complexity, we select the chirp filtering technique to compensate for CD. Then, we configure a scanning delay algorithm(SDA) to estimate the TX IQ skew. The SDA estimates and compensates the TX IQ skew by finding the value that yields the lowest equalizer error. Second, we propose a low-complexity equalizer to compensate for the TX amplitude and phase imbalance. We combine a complex-valued MIMO(CV-MIMO) and a TMMA-RV-MIMO for signal equalization, which employs a butterfly and a non-butterfly structure, respectively. The CV-MIMO is used to perform polarization demultiplexing. The TMMA-RV-MIMO equalizes each of the two polarizations and simultaneously compensates for the TX IQ imbalance. Since the TX and RX skew are being compensated, the new equalizer can equalize the signal with fewer taps. For convenience, the new equalizer is called C-R AEQ. The 36 GBaud DP-16QAM simulation and experiment results demonstrate that the RX IQ skew estimation error is less than 0.25 ps, and the TX IQ skew estimation error is less than 0.9 ps. The C-R AEQ tolerates phase imbalance up to ±5 degrees and amplitude imbalance up to ±0.1 with a 0.3 dB Q-factor penalty. The number of real multiplications is reduced by 50% compared with conventional cases in total.
@@ 2.1 Working principle
Where the is , and is . According to Eq. 2, the delay introduces respective linear phases shift to and in frequency domain. Therefore, the transfer function of delay in the frequency domain can be calculated as
@@
However, if there are IQ skews in X polarization or Y polarization, the numerically controlled oscillator(NCO) outputs leading to the error sampling instant, as shown in Fig. 1(b). So, if we calculate the I and Q separately, as shown in Fig.1(c), the IQ skew can be determined as
Where , are the receiver IQ skew of the X polarization and Y polarization(pol), Baud is the symbol rate. 2Baud is the sample rate. The notation mean() represents the average calculation. are fractional intervals of I and Q in each Pol.
Fig.2 illustrates Gardner’s skew monitoring algorithm(GSMA), deployed at the receiver, which employs Gardner’s PD as its timing error detector(TED). The algorithm inputs the I and Q components of both polarizations (Xi, Xq, Yi, Yq) into four separate CRAs. The outputs of the GSMA are the fractional interval , and . Using these values, the IQ skew for the corresponding polarization is calculated via Eq.5 and Eq.6.
However, the effectiveness of the GSMA is impaired by CD compensation. When CD is compensated at the receiver, the algorithm fails to accurately estimate the RX IQ skew. This limitation is analyzed using the WL model in [4]. Let ,we can write . Thus, the , which suffers from CD and receiver IQ skew, can be represented as
@@
Where is the signal after CD compensation in the frequency domain. The GSMA require estimate the phase shift from . However, as seen from Eq.(12), does not maintain a strictly linear phase. Consequently, the receiver IQ skew becomes undetectable when CD compensation is applied at the receiver.
This difference is illustrated in Fig.(3). Fig. (3)(a) shows the phase shift between the delayed and original signals in the frequency domain. The delayed signal always keeps a fixed phase difference between the delayed signal and the original signal. Fig.3(b) shows the phase shift between the and the original signal in the frequency domain. It is observed in Fig.(3)(b) that the phase difference changes when the changes. Therefore, the delay cannot be detected by the GSMA.
The IQ pha./amp imbalance is mixed in the same way. Assuming the signal is added amplitude and phase imbalance, which can be expressed as
Where g is the amplitude imbalance, and is the phase imbalance. If the signal suffers from CD, which can be expressed as
Similar to Eq. (12), the is used to compensate for CD, and the result can be expressed as
contains a component that consists of the complex conjugate of the signal. This component acts as interference that degrades the performance of the equalizer. According to Eq. (15), it can be observed that the impact of this conjugate term on is lower when the and g are smaller. Thus, the equalizer can partly compensate for the IQ pha./amp. imbalance when these imbalances are small.
Based on the analysis above, we propose a DSP flow as shown in Fig. 4. In TX DSP, to address the issue that the receiver’s IQ skew cannot be extracted under large CD, we implement the CD compensation at the TX side. Furthermore, to reduce the complexity of CD compensation, we employ the chirp filter technique. The transfer function of the pre-chirp filter can be expressed as[22]
@@
The scanning delay algorithm(SDA) is employed to compensate for the TX IQ skew. Figure 5 is the diagram of the SDA, which comprises a scanning delay block, a scan controller block, and a TX IQ skew compensation block. Within the scanning delay block, the X pol and Y pol signals are subject to a gradually increasing delay from to , where is an integer. This process can be expressed as
Where the or represent the IQ skew of X pol or Y pol, respectively. The and is gradually increasing delay from to . In the RX DSP, the CV-MIMO module outputs the to the scan controller block after the scanning delay block. Subsequently, the scan controller block calculates the error . The can be expressed as
Where N is the number of symbols, / is X/Y pol equalization error. During the scanning process, the scan controller block records a set of , and identifies the minimum . Then, the and that correspond to this minimum are output as the estimated TX skew. Subsequently, the TX IQ skew compensation block uses these estimated delays to compensate for the TX IQ skew. The RX IQ pha./amp. imbalance is compensated for using GSOP. The RX IQ skew is compensated for using GSMA. The GSMA diagram is shown in Fig. 2.
To compensate for the TX IQ amp./pha. imbalance and equalize signal with low complexity, we propose a C-R AEQ inspired by the KM AEQ [24]. The C-R AEQ is composed of CV-MIMO and TMMA-RV-MIMO. The CV-MIMO is composed of a butterfly structure with fewer taps per filter, as we found that frequency offset estimation (FOE) can be performed correctly after the signal is equalised by fewer taps CV-MIMO(such as 3-tap). The TMMA-RV-MIMO is a non-butterfly filter after the FOE as shown in Fig. (8). The CV-MIMO is used for polarization demultiplexing and to ensure that FOE is correctly performed. TMMA-RV-MIMO equalizes each of the two polarisations and compensates for the TX IQ imbalance. Moreover, due to the phase sensitivity of TMMA-RV-MIMO, the TMMA-RV-MIMO can recover the carrier phase(CPR). The tap coefficient vector of the TMMA-RV-MIMO is updated according to ref. [1]. C-R AEQ is similar to ref.[23]. The difference is that ref.[23] does not consider CD and frequency offset. This means ref.[23] is only suitable for a self-homodyne coherent optical system in a short transmission length. And DD-LMS requires symbol decision , whereas C-R AEQ does not require it. In addition, ref.[23] requires higher complexity.